# Behavioral Refinement via Interpolant-based Policy Diffusion

### v1

@@ Appendix B Experiment Domains
Adroit (State). Adroit encompasses the control of a 24-degree-of-freedom robotic hand to accomplish four specific tasks (Fig. 6). Adroit is one of the most challenging task sets, as it requires a model to generate intricate, high-dimensional, and high-precision action sequences over long time horizons. We use the dataset of human demonstrations provided in the DAPG repository [13]. Similar to Franka Kitchen, we train the model on different data sizes: 1) small size: $1250$ sequences, 2) medium size: $2500$ sequences, and 3) large size: $5000$ sequences. We evaluate two source distributions: 1) CVAE [44] trained on the same dataset (on $4$ tasks) and 2) hand-craft heuristic source distribution (on door and hammer). The models are trained on state observation. Similar to Franka Kitchen, we apply receding-horizon control; the models predict $64$ steps of actions, of which $48$ steps of actions are executed on the robot without re-planning.
6-DoF Grasp Pose Generation (Point-Cloud). The objective is to generate grasp poses capable of picking up an object based on the object’s point cloud (Fig. 6). The demonstrated behavior’s distribution is intricate and highly multi-modal [47], with high-dimensional point clouds as observations. Our model is trained on $11$ object categories sourced from the Acronym dataset [48], encompassing $552$ objects with approximately $200$-$2000$ grasps for each object. Similar to Adroit, we evaluate two different source distributions: 1) CVAE [44] trained on the same dataset and 2) hand-crafted heuristic source distributions. Both source distributions use point-cloud observations.
@@ Appendix C Source Policies
Data-driven policies. We use lightweight Conditional Variational Autoencoders (CVAEs) as a representative policy. The CVAEs are trained to generate action sequences conditioned on observations specific to each domain (as described above). In Franka Kitchen and Adroit tasks, the CVAEs are parameterized by multi-layer perceptrons with $\approx 3\times 10^{6}$ parameters ($10$ times fewer parameters compared to the diffusion models). In Grasp Generation task, CVAE is parameterized using the Grasp SE(3)-DiffusionFields neural network [47] with $\approx 3.5\times 10^{6}$ parameters (approximately half of the number of parameters of the diffusion models).
Table III: Values of $\gamma$ scale $d$ and $\epsilon$ scale $c$ in BRIDGeR under different source policies.
@@ Appendix D Baselines
We compare BRIDGeR Policy with the current state-of-the-art diffusion-style behavior cloning. Since BRIDGeR can be viewed as one kind of residual model, we also compare it against a residual policy. For ablation studies, we evaluate BRIDGeR under different configurations, focusing on the source distributions and interpolants. In the following, we present an overview of each methods.
DDIM. Diffusion Policy [6, 36] trains a Denoising Diffusion Probabilistic Model [16] and applied Denoising Diffusion Implicit Model (DDIM) [45] during testing. In Adroit and Franka Kitchen tasks, the model is parameterized by the CNN-based U-net [6]. In Grasp Generation tasks, the model is parameterized by Grasp SE(3)-DiffusionFields neural network [47].
@@
BRIDGeR. Our stochastic interpolant policy. The model is parameterized by the CNN-based U-net [6] for the Adroit and Franka Kitchen tasks and parameterized by Grasp SE(3)-DiffusionFields neural network [47] for Grasp Generation. When compared with other baselines, we set $\gamma$ scale $d$ and $\epsilon$ scale $c$ as in Table. III. We conduct an ablation study of BRIDGeR with different source distribution, interpolant functions, $\gamma$ and $\epsilon$ scale. We name each ablation as BRIDGeR / []delimited-[][$\cdots$][ ⋯ ] / []delimited-[][$\cdots$][ ⋯ ]. For example, BRIDGeR / Linear / CVAE represents BRIDGeR with CVAE as the source distribution and linear function as the interpolant.
For fair comparison, each method has approximately the same number of parameters. In Adroit and Franka Kitchen, methods have $\approx 6.7\times 10^{7}$ parameters each. In Grasp Generation tasks, the parameter count for each method is $\approx 6.2\times 10^{6}$.
@@ Appendix E Model Training
In both Adroit and Franka Kitchen, we utilized AdamW [25] as the optimizer. The model was trained for $3000$ epochs for Adroit and $5000$ epochs for Franka Kitchen. The learning rate was set to $5\times 10^{-6}$ with a learning rate decay of $0.5$ every $500$ epochs. The batch size was fixed at $256$.
@@ Appendix F Additional Results and Plots
BRIDGeR against baselines. Performance comparisons between BRIDGeR v.s. the baselines on different training data sizes and diffusion steps for Adroit and Franka Kitchen are shown in Table. IV (large training data size), V (medium training data size) and VI (small training data size). Results for Grasp Generation with varying number of steps are presented in Table. VII and VIII. We report the results of BRIDGeR with Power3 interpolant function and decomposed velocity $b$. The values of $\gamma$ scale and $\epsilon$ scale are listed in Table. III.
BRIDGeR with Power3 and Linear interpolant function. The performance metrics comparing the different interpolant functions on the Adroit and Franka Kitchen tasks are shown in Table. IX (large training data size), X (medium training data size) and XI (small training data size). The results for grasp generation presented in Table. XII and XIII.
F-A Other Findings
@@
Table IV: Full results of average task performance of BRIDGeR against current state-of-the-art methods under varying number of diffusion steps when trained with the Large dataset.
@@
Table V: Full results of average task performance of BRIDGeR against current state-of-the-art methods under varying number of diffusion steps when trained with the Medium dataset.
@@
Table VI: Full results of average task performance of BRIDGeR against current state-of-the-art methods under varying number of diffusion steps when trained with the Small dataset.
@@
Table VII: Success rate (averaged over $100$ grasps on ten test objects). BRIDGeR significantly outperforms DDIM and Residual Policy across the number of diffusion steps. Compared to SE3, BRIDGeR achieve higher success rate when the number of diffusion steps is small. We show up to $k=160$ steps to be consistent with prior reported results [47].
@@
Table VIII: Average Earth Mover’s Distance between generated grasp poses and target grasp poses. BRIDGeR significantly outperforms SE3, DDIM and Residual Policy across the number of diffusion steps..
@@
Table IX: Full results of average task performance of BRIDGeR with Power3 and Linear interpolant functions under varying number of diffusion steps when trained with the Large dataset.
@@
Table X: Full results of average task performance of BRIDGeR with Power3 and Linear interpolant functions under varying number of diffusion steps when trained with the Medium dataset.
@@
Table XI: Full results of average task performance of BRIDGeR with Power3 and Linear interpolant functions under varying number of diffusion steps when trained with the Small dataset.
@@
Table XII: Success rate (averaged over $100$ grasps on ten test objects). BRIDGeR with Power3 interpolant function performs better than Linear interpolant function. We show up to $k=160$ steps to be consistent with prior reported results [47].
@@
Table XIII: Average Earth Mover’s Distance between generated grasp poses and target grasp poses. BRIDGeR with Power3 interpolant function performs better than Linear interpolant function.

### v2

@@ Appendix B Experiment Domains
Adroit (State). Adroit encompasses the control of a 24-degree-of-freedom robotic hand to accomplish four specific tasks (Fig. 6). Adroit is one of the most challenging task sets, as it requires a model to generate intricate, high-dimensional, and high-precision action sequences over long time horizons. We use the dataset of human demonstrations provided in the DAPG repository [14]. We evaluate a model’s performance using the proportion of task successes. An episode is considered successful if the cumulative reward exceeds a threshold ($9$ for Door and $10$ for the rest of the tasks). This threshold is selected to ensure task completion. For instance, in the case of the Adroit Door task, a reward greater than $9$ is received if the door is successfully opened. Similar to Franka Kitchen, we train the model on different data sizes: 1) small size: $1250$ sequences, 2) medium size: $2500$ sequences, and 3) large size: $5000$ sequences. We evaluate two source distributions: 1) CVAE [45] trained on the same dataset (on $4$ tasks) and 2) hand-craft heuristic source distribution (on door and hammer). The models are trained on state observation. Similar to Franka Kitchen, we apply receding-horizon control; the models predict $64$ steps of actions, of which $48$ steps of actions are executed on the robot without re-planning.
6-DoF Grasp Pose Generation (Point-Cloud). The objective is to generate grasp poses capable of picking up an object based on the object’s point cloud (Fig. 6). The demonstrated behavior’s distribution is intricate and highly multi-modal [49], with high-dimensional point clouds as observations. Our model is trained on $11$ object categories sourced from the Acronym dataset [50], encompassing $552$ objects with approximately $200$-$2000$ grasps for each object. Similar to Adroit, we evaluate two different source distributions: 1) CVAE [45] trained on the same dataset and 2) hand-crafted heuristic source distributions. Both source distributions use point-cloud observations. For real-world evaluation, we use the models trained in simulated grasp generation and test it on $10$ real world objects (Fig. 13). For each object, we evaluated $10$ grasps uniformly sampled from a set of filtered grasps. Each model generated $2000$ candidate grasps, which we filtered using hand-crated rules (e.g., we remove unreachable grasps). The ratio of filtered grasps were similar among different models (BRIDGeR / CVAE: $57.52\%$ vs. BRIDGeR / Heuristic: $58.12\%$ vs. SE3: $55.03\%$)
Cleaning (Image and State). The robot had to grasp a sponge and wipe marks off a surgical practice model that mimics human tissue. The demonstrated behavior’s distribution is high-dimensional (22 dimensions of robot joint angles per time-step), multi-modal demonstrations (e.g., wiping the same mark either from left to right or from up to down), and manipulating a deformable sponge. Our model was trained on $60$ collected demonstrations. Each trajectory involves grasping the sponge and clean the markers which are randomly positioned on the surgical sample. Similar to Adroit, we evaluate two different source distributions: 1) CVAE trained on the same dataset and 2) hand-crafted heuristic source distributions. We apply receding-horizon control; the models predict $48$ steps of actions, of which $16$ steps of actions are executed on the robot without re-planning. We evalute the model performance by normalized cleaned area, which is calculated by the ratio of the area of the cleaned marker to the original area of the marker. (Fig. 15).
Figure 15: The original view of human tissue is transformed by perspective wrap. After transformation, we calculate the normalized cleaned area by the ratio of the cleaned area (green dashed border) to the original area of the marker (red dashed border). The marker and cleaned area are segmented by thresholding the pixel values.
@@ Appendix C Source Policies
  • Cleaning: We recorded a single fixed reference trajectory, which moves the robot arm from an initial pose to grasp the sponge and cleans a specific position on the surgical model. To sample from source policy, we compute the closest robot state (joint angles) in the reference trajectory to the current robot state, extract the subsequent $48$ actions, and add Gaussian noise. The methods must learn to grasp the sponge from various initial poses and clean marks positioned at different locations.
Data-driven policies. We use lightweight Conditional Variational Autoencoders (CVAEs) as representative policies. The CVAEs are trained to generate action sequences conditioned on observations specific to each domain. In Franka Kitchen and Adroit, the CVAEs are parameterized by multi-layer perceptrons with $\approx 3\times 10^{6}$ parameters ($10$ times fewer parameters compared to the diffusion models). In Grasp Generation, CVAE is parameterized using the Grasp SE(3)-DiffusionFields neural network [49] with $\approx 3.5\times 10^{6}$ parameters (approximately half of the number of parameters of the diffusion models). In Cleaning, the CVAE is parameterized by a multi-layer perceptron with $\approx 3\times 10^{6}$ parameters, along with image and state encoders with $\approx 1\times 10^{7}$ parameters.
Table VI: Values of $\gamma$ scale $d$ and $\epsilon$ scale $c$ in BRIDGeR under different source policies.
@@ Appendix D Baselines
We compare BRIDGeR Policy with the current state-of-the-art diffusion-style behavior cloning. Since BRIDGeR can be viewed as a type of (probabilistic) residual model, we also compare it against a residual policy. For ablation studies, we evaluate BRIDGeR under different configurations, focusing on the source distributions and interpolants. In the following, we present an overview of each methods.
DDIM. Diffusion Policy [7, 37] trains a Denoising Diffusion Probabilistic Model [17] and applied Denoising Diffusion Implicit Model (DDIM) [46] during testing. In Adroit and Franka Kitchen tasks, the model is parameterized by the CNN-based U-net [7]. In Grasp Generation tasks, the model is parameterized by Grasp SE(3)-Diffusion Fields neural network [49]. In Cleaning, the model is parameterized by the CNN-based U-net [7] similar to the Adroit and Franka Kitchen tasks, but with an additional image and state encoders [7].
@@
BRIDGeR. Our stochastic interpolant policy. The model is parameterized by the CNN-based U-net [7] for the Adroit, Franka Kitchen and Cleaning (with an additional image and state encoders similar to DDIM) tasks and parameterized by Grasp SE(3)-DiffusionFields neural network [49] for Grasp Generation. When compared with other baselines, we set $\gamma$ scale $d$ and $\epsilon$ scale $c$ as in Table. VI.
For fair comparison, each method has approximately the same number of parameters. In Adroit and Franka Kitchen, methods have $\approx 6.7\times 10^{7}$ parameters each. In Grasp Generation, the parameter count for each method is $\approx 6.2\times 10^{6}$ and in Cleaning, the number of parameters is $3.0\times 10^{7}$ for each method.
@@ Appendix E Model Training
In both Adroit, Wound Cleaning and Franka Kitchen, we utilized AdamW [26] as the optimizer. The model was trained for $3000$ epochs for Adroit and Wound Cleaning, and $5000$ epochs for Franka Kitchen. The learning rate was set to $5\times 10^{-6}$ with a learning rate decay of $0.5$ every $500$ epochs. The batch size was fixed at $256$.
@@ Appendix F Additional Results and Plots
BRIDGeR against baselines. Performance comparisons between BRIDGeR v.s. the baselines on different training data sizes and diffusion steps for Adroit and Franka Kitchen are shown in Table. VII (large training data size), VIII (medium training data size) and IX (small training data size). Results for Grasp Generation with varying number of steps are presented in Table. X and XI. We report the results of BRIDGeR with Power3 interpolant function and decomposed velocity $b$. The values of $\gamma$ scale and $\epsilon$ scale are listed in Table. VI.
BRIDGeR with Power3 and Linear interpolant function. The performance metrics comparing the different interpolant functions on the Adroit and Franka Kitchen tasks are shown in Table. XII (large training data size), XIII (medium training data size) and XIV (small training data size). The results for grasp generation presented in Table. XV and XVI.
@@
Table VII: Full results of average task performance of BRIDGeR against current state-of-the-art methods under varying number of diffusion steps when trained with the Large dataset.
@@
Table VIII: Full results of average task performance of BRIDGeR against current state-of-the-art methods under varying number of diffusion steps when trained with the Medium dataset.
@@
Table IX: Full results of average task performance of BRIDGeR against current state-of-the-art methods under varying number of diffusion steps when trained with the Small dataset.
@@
Table X: Success rate (averaged over $100$ grasps on ten test objects). BRIDGeR significantly outperforms DDIM and Residual Policy across the number of diffusion steps. Compared to SE3, BRIDGeR achieve higher success rate when the number of diffusion steps is small. We show up to $k=160$ steps to be consistent with prior reported results [49].
@@
Table XI: Average Earth Mover’s Distance between generated grasp poses and target grasp poses. BRIDGeR significantly outperforms SE3, DDIM and Residual Policy across the number of diffusion steps..
@@
Table XII: Full results of average task performance of BRIDGeR with Power3 and Linear interpolant functions under varying number of diffusion steps when trained with the Large dataset.
@@
Table XIII: Full results of average task performance of BRIDGeR with Power3 and Linear interpolant functions under varying number of diffusion steps when trained with the Medium dataset.
@@
Table XIV: Full results of average task performance of BRIDGeR with Power3 and Linear interpolant functions under varying number of diffusion steps when trained with the Small dataset.
@@
Table XV: Success rate (averaged over $100$ grasps on ten test objects). BRIDGeR with Power3 interpolant function performs better than Linear interpolant function. We show up to $k=160$ steps to be consistent with prior reported results [49].
@@
Table XVI: Average Earth Mover’s Distance between generated grasp poses and target grasp poses. BRIDGeR with Power3 interpolant function performs better than Linear interpolant function.